5. Assume a random sample of the birth weights of 186 healthy babies has a mean of 3103 g and ( pomt) a standard deviation of 66 g Construct a 95% confidence-terval estimate of the mean weight of all healthy babies born to healthy mothers. What does the interval suggest about a study informing soon-to-be-parents that they can expect their new baby to weigh about 2980 g 03002 g <pc3204 g·the 2980 g weght s less than the range values of enterval, however, tis close to he menal and could therefre easily represent an expected mean weight of ew babes 03002 g 〈パ3204 the NSO g weight falls below he mterval, suggest ng that it sun kely to be an expected value and s l kely too low to repres ntanaccurate popu ation mean. 2879gA3080 g, the 2980 g weight falls comfortably in the interval, which suggests that 2980 g is very likely to repeesent an accurate population mean of birth weights. 02579 g<バ3080 g, the interval suggests 3103 gsnot a likely mean proving the sample s based, the recent study should be used as a bass rest ating tw be he ghts. Assume you wat to construct a 98% confidence merval with a sample ofr-lo tom ofa omally distributed population. Find the critical value g2 6 dpotro 0 2764 02.821 01.383 3.250 7. A sample of size -12 does sot have a known population standard deviation The populatice ( point appears to be normally distnibuted. Determine whether a margin of eror should be calculated using a critical value of ig2, a critical value of tg2, or neither O a critical value of fa 2 O a critical value of ig2 neither
8·The standard deviation of sample size of n-15 is 345 and the mean is 6784. Given that the (1 point) population follows a normal distribution, construct a 90% confidence interval estimate of the mean of the population 06519 < 7049 6593 < μ < 6975 6627 < μ < 6941 6783 < < 6785 9. You are constructing a 90% confidence interval for a sample consisting of n = 9 values and an (1 point) unknown population standard deviation. The population appears to be very skewed. Determine whether a margin of error should be calculated using a critical value of za2, a critical value of a2, or neither a critical value of a2 O a critical value of ta2 O neither 10. A simple random sample has a sample size of n- 65. Given the population is normally (1 potnt) distributed, find the critical value ta2 corresponding to a 99% confidence level. 2.678 02.575 2.000 2.660
Answer #1

5)

Level of Significance ,    α =    0.05
sample std dev ,    s =    696
Sample Size ,   n =    186
Sample Mean,    x̅ =   3103

degree of freedom=   DF=n-1=   185
't value='   tα/2= t(0.025,185) =   1.9729
      
Standard Error , SE =   s/√n =   51.0332
margin of error ,   E=t*SE =   100.6819
confidence interval is       
Interval Lower Limit=   x̅ - E =    3002.3181 (3002g)
Interval Upper Limit=   x̅ + E =    3203.6819(3204g)

Answer: option b) is answer
-----------------------------------------------------------------------------------------------------------------

only one ques per post can be answered as HOMEWORKLIB RULES answering guidelines